OpenStudy (anonymous):

Solve: 3/4m-1/3=1/2 Not answer just steps

OpenStudy (anonymous):

\[-\frac{1}{3}+\frac{3 m}{4}=\frac{1}{2} \]Multiply both sides by a number that 3, 4 and 2 will divide into evenly. 12 will do.\[12\left(-\frac{1}{3}+\frac{3 m}{4}\right)==\left(\frac{1}{2}\right)12 \]Simplified:\[-4+9 m=6 \]Put -4 on the right side and divide the result by 9.\[m=\frac{(6+4)}{9}=\frac{10}{9} \]To check the result, set m to 10/9.\[-\frac{1}{3}+\frac{3 (10/9)}{4}=\frac{1}{2} \]